# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v2

@@ 3.3.1. Workflow architecture
Communicative architectures can facilitate debate and conversation among workers (Du et al., 2023; Teevan et al., 2016; Wu et al., 2023a; Salisbury et al., 2017; Drapeau et al., 2016; Chen et al., 2019; Chang et al., 2017; Kobayashi et al., 2018; Liang et al., 2023) and allow workers to return feedback to the user (Kulkarni et al., 2012; Lasecki et al., 2015; Nebeling et al., 2016). Some workflows create roles for different workers, often based on roles held in real-life organizations (e.g., journalists, software development teams, etc.) (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014; Park et al., 2022). Some papers integrate these roles into a hierarchy of communication, with managers, midlevel managers, and different forms of communication between different levels of the hierarchy (Agapie et al., 2015; Retelny et al., 2014). Communication may be restrained through subtasks, free-form though email (Agapie et al., 2015; Valentine et al., 2017), or enabled through a shared interface (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013). Communication among workers may be synchronous or asynchronous. Synchronous workflows help enable collaboration (André et al., 2014; Chilton et al., 2014; Salisbury et al., 2017; Chilton et al., 2019) and can reduce latency (Bernstein et al., 2011). With asynchronous workflows or workflows with limited communication architectures, a working memory or state management can help workers access progress reports, what other workers think are good next steps, and the overarching goals (Valentine et al., 2017; Hahn et al., 2016; Lasecki et al., 2013; LaToza et al., 2014; Zhang et al., 2012; Hahn et al., 2016; Yang et al., 2022; Nair et al., 2023b; Wu et al., 2023a). Communicative architectures can serve as a method of disagreement resolution and workflow flexibility. They often encourage diverse responses though worker communication.
3.3.2. Subtasks
Generate. Most workflows involve at least one step in which workers generate content. This generation may be writing words or code (Pavel et al., 2014; Weir et al., 2015; Zhang et al., 2022; Li et al., 2023; Zhang et al., 2023a), answering a question or solving a problem (Kobayashi et al., 2018; Wu et al., 2023b; Xie et al., 2023; Eisenstein et al., 2022), or brainstorming and ideation (Chilton et al., 2019; Lasecki et al., 2013; Mahyar et al., 2018; Huang et al., 2018; Long et al., 2023; Yao et al., 2023). Generation tasks may also ask for labels or annotations on given content (Guo et al., 2017, 2019; Chilton et al., 2014; Chang et al., 2017; Bragg et al., 2013), to translate between languages (Ambati et al., 2012; Zaidan and Callison-Burch, 2011), to contribute to a conversation with other workers (Mahyar et al., 2018; Wang et al., 2022; Liang et al., 2023; Du et al., 2023), to demonstrate a behavior (Bragg et al., 2021; Lee et al., 2017), or to source content from external sources (Kim et al., 2014; Chilton et al., 2019; Hahn et al., 2016; Kim and Monroy-Hernandez, 2016; Pietrowicz et al., 2013; Huang et al., 2015; Luther et al., 2015; Christoforaki and Ipeirotis, 2014). Some subtasks also require an explanation along with the generated content (Chang et al., 2017; Kim et al., 2017; Cheng and Bernstein, 2015; Drapeau et al., 2016; Schick et al., 2022; Gero et al., 2023; Fleming et al., 2023).
Evaluate. One of the most common evaluation tasks is choosing the best among multiple possible options (Guo et al., 2017; Weir et al., 2015; Guo et al., 2019; Cheng and Bernstein, 2015; Kittur et al., 2012; Liu et al., 2018). Other subtasks may include rating and ranking among multiple options (Pavel et al., 2014; Salehi et al., 2017; Marcus et al., 2011a; Huffaker et al., 2020; Willett et al., 2012; Mahyar et al., 2018), a pairwise comparison (Mohanty et al., 2019; Little et al., 2010b, a; Marcus et al., 2011a, b; Parameswaran et al., 2023; Bursztyn et al., 2022). Evaluation subtasks may also determine if another round of generation needs to occur (Wu et al., 2022b; LaToza et al., 2014; Wang et al., 2022; Liang et al., 2023).
Workflows can decide what moves forward through majority vote (Huang et al., 2017; Dai et al., 2010; Lasecki et al., 2015; Liu et al., 2018; Kim et al., 2014; Guo et al., 2019), other aggregations of voting tasks (e.g., expectation maximization (Drapeau et al., 2016; Parameswaran et al., 2023), weighted mean (Mohanty et al., 2019), weak supervision (Arora et al., 2022)), returning the highest ranked or rated items (Pavel et al., 2014; Huffaker et al., 2020), or automated methods. These automated methods include measuring agreement from the generation step (Liu et al., 2018; Hahn et al., 2016; Huffaker et al., 2020; Chang et al., 2017; Huang et al., 2015; Nair et al., 2023b; Bernstein et al., 2011) and using non-AI checks like semantic embeddings and rule-based heuristics (Parameswaran et al., 2023; Wang et al., 2023b; Ambati et al., 2012). There are also subtasks that seek critiques and reviews from the user (Nebeling et al., 2016; Schick et al., 2022) or another worker (Kim et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Christoforaki and Ipeirotis, 2014; Mahyar et al., 2018; Schick et al., 2022; Parameswaran et al., 2023; Madaan et al., 2023b; Liang et al., 2023). These subtasks may ask for feedback or verify if errors or other qualities exist (Huang et al., 2023b; Bragg et al., 2013; Christoforaki and Ipeirotis, 2014; Chilton et al., 2019; Weir et al., 2015; Huang et al., 2015; Kim et al., 2014; Gero et al., 2023; Huang et al., 2017; Bernstein et al., 2010; Yang et al., 2022; Wu et al., 2023c; Kittur et al., 2011; Liang et al., 2023; Kulkarni et al., 2012; André et al., 2013; Butler et al., 2017). The outputs from these subtasks can disqualify information or pass into another subtask to be addressed. Evaluate subtasks primarily resolve disagreements and serve the verification step strategy.
Transform. Other subtasks transform given data by merging or improving it. Some subtasks merge multiple streams of incoming content. This merge could be algorithmic (Alshaibani et al., 2020; Ahmad et al., 2011; Alshaibani and Quinn, 2021; Teevan et al., 2016; Chilton et al., 2019; Kittur et al., 2011) or accomplished by the worker (Kittur et al., 2011; Wu et al., 2023c, 2022b; Lee et al., 2017; LaToza et al., 2014; Ambati et al., 2012; Teevan et al., 2016; Kittur et al., 2012; Wu et al., 2023c; Pietrowicz et al., 2013; Luther et al., 2015; Kulkarni et al., 2012). Other subtasks improve content generated by previous steps, sometimes by incorporating feedback from validation steps. Improving on previously generated content can come in many forms: updating the format of the content (Huang et al., 2023b; Wu et al., 2022b; Eisenstein et al., 2022; Wang et al., 2022; Arora et al., 2022), pruning bad results (Liu et al., 2018; Gero et al., 2023), adding missing results (Huang et al., 2023b; Liu et al., 2018; Weir et al., 2015), correcting errors (Huang et al., 2023b; Yang et al., 2022; Drapeau et al., 2016; Kobayashi et al., 2018; Gouravajhala et al., 2018), and performing direct improvements (Wu et al., 2023c; Madaan et al., 2023b; Hahn et al., 2016; Mahyar et al., 2018; Lee et al., 2017; Huffaker et al., 2020; Nebeling et al., 2016; Christoforaki and Ipeirotis, 2014; Little et al., 2010b; Liem et al., 2011; Dai et al., 2010; Schick et al., 2022; Kim et al., 2017; Bernstein et al., 2010; Guo et al., 2019; Teevan et al., 2016; McDonnell et al., 2016). Transform subtasks primarily support the strategies of resolving disagreements among diverse responses.
Focus. One set of subtasks looks to select the attention of the upcoming tasks within a broader context. For example, this step may find patches of the input that need editing, extract information from context, or focus on common keywords (Bernstein et al., 2011; Huffaker et al., 2020; Gero et al., 2023; Creswell and Shanahan, 2022; Creswell et al., 2022; Nair et al., 2023b; Kim et al., 2017). Focus subtasks address the strategies of workflow flexibility, worker instructions for other subtasks, and creating appropriate subtasks.
Partition. Some subtasks transform a task by breaking it up into smaller pieces (Kittur et al., 2011; LaToza et al., 2014; Kulkarni et al., 2012; Wu et al., 2023c; Zhou et al., 2022b; Valentine et al., 2017; Wu et al., 2022b; Huang et al., 2023c). For example, the Divide step in the Turkomatic workflow asks workers to divide a given subtask into simpler subtasks. The workflow calls this step recursively until the subtask complexity is valued (by a separate crowdworker) at a specified cost (Kulkarni et al., 2012). Partition subtasks create appropriate subtasks and support workflow flexibility.
3.3.3. Workers
Workers that complete the subtasks can be crowdworkers, LLMs, or the user.
Crowdworkers are usually recruited from crowdworking platforms like Amazon Mechanical Turk or UpWork (formerly known as oDesk). Platforms allow for requests of worker qualifications, although these credentials are sometimes hard to validate (Ahmad et al., 2011; Zaidan and Callison-Burch, 2011; Valentine et al., 2017; Kulkarni et al., 2014). Methods like retainer models can reduce the latency of accessing crowdworkers on these platforms (Bernstein et al., 2011). Workers can also be recruited from the public, including community members (Mahyar et al., 2018), conference committees (Chilton et al., 2014; André et al., 2013), and online deaf communities (Bragg et al., 2021).
Crowdworkers have limitations that overlap with LLM limitations. Prior work suggests that approximately of outcomes from Amazon Mechanical Turk are low quality (Bernstein et al., 2010), and may include spam or off-topic responses (Butler et al., 2017; Lasecki et al., 2013; Alshaibani et al., 2020; Huang et al., 2018). Crowdworkers also struggle with following long and complex instructions, as people can overly focus on the beginning or end of a given text (Hahn et al., 2016; Chilton et al., 2013; Lasecki et al., 2015; Chen et al., 2019). Crowdworkers can be biased; for example, in one study, workers performed significantly better in identifying photos of their own race compared to other races (Mohanty et al., 2019).
LLMs. There are many possible LLMs of different sizes and expenses (Bursztyn et al., 2022). Models have differing strengths and respond differently to prompts (Mirowski et al., 2023; Wu et al., 2023c; Parameswaran et al., 2023; Liang et al., 2023). LLM outputs can be adjusted through parameter changes (e.g., temperature) (Park et al., 2022; Parameswaran et al., 2023), fine-tuning (Creswell et al., 2022; Creswell and Shanahan, 2022; Cobbe et al., 2021; Li et al., 2023; Bursztyn et al., 2022), and expressing roles through prompts (Park et al., 2022; Du et al., 2023).
Some LLM limitations overlap with those of crowdworkers. LLMs can provide erroneous outputs: they are known to confidently output incorrect answers (Parameswaran et al., 2023; Wang et al., 2023b; Du et al., 2023) or hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and for struggling with self-consistency and logical reasoning (Creswell et al., 2022; Creswell and Shanahan, 2022; Yao et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Xie et al., 2023). LLMs can ignore large fractions of long prompts or focus mostly on recent parts of the given context (Parameswaran et al., 2023; Du et al., 2023; Wu et al., 2023c; Madaan et al., 2023b). LLMs also output biased results, a widespread concern (Park et al., 2022; Wu et al., 2023a; Huang et al., 2023a; Long et al., 2023). For example, work on screenplay generation found predictable and stereotypical outputs, with multiple instances of gender biases (Mirowski et al., 2023). Furthermore, LLM training on human data increases the risk of accidental plagiarism (Mirowski et al., 2023).
[…]

### v3

@@ 4.2.3. Workflow architecture
Communicative architectures can facilitate debate and conversation among workers (Du et al., 2023; Teevan et al., 2016; Wu et al., 2023a; Salisbury et al., 2017; Drapeau et al., 2016; Chen et al., 2019; Chang et al., 2017; Kobayashi et al., 2018; Liang et al., 2023) and allow workers to return feedback to the user (Kulkarni et al., 2012; Lasecki et al., 2015; Nebeling et al., 2016). Some workflows create roles for different workers, often based on roles held in real-life organizations (e.g., journalists, software development teams, etc.) (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014; Park et al., 2022). Some papers integrate these roles into a hierarchy of communication, with managers, midlevel managers, and different forms of communication between different levels of the hierarchy (Agapie et al., 2015; Retelny et al., 2014). Communication may be restrained through subtasks (Chang et al., 2017; Chen et al., 2019), free-form though email (Agapie et al., 2015; Valentine et al., 2017), or enabled through a shared interface (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013). Communication among workers may be synchronous or asynchronous. Synchronous workflows help enable collaboration (André et al., 2014; Chilton et al., 2014; Salisbury et al., 2017; Chilton et al., 2019) and can reduce latency (Bernstein et al., 2011). With asynchronous workflows or workflows with limited communication architectures, a working memory or state management can help workers access progress reports, what other workers think are good next steps, and the overarching goals (Valentine et al., 2017; Hahn et al., 2016; Lasecki et al., 2013; LaToza et al., 2014; Zhang et al., 2012; Hahn et al., 2016; Yang et al., 2022; Nair et al., 2023b; Wu et al., 2023a).
5. Strategies integrating the design space
Choosing tactic designs to support objectives is a non-trivial task, as combining different tactic types and implementations supports a huge range of possible workflows. However, through the thematic coding process we surfaced strategies by which tactics are used to support objectives (Figure 4). Specifically, these strategies (5.1) encourage diverse responses, (5.2) validate outputs from subtasks, (5.3) enable user guidance, (5.4) utilize adaptable architectures, (5.5) improve worker recruitment, (5.6) align a subtask’s required skills with a worker’s capabilities, (5.7) promote global context, and (5.8) ensure outputs meet a predetermined quality thresholds.
We outline the strategies that surfaced from the literature in a manner adapted from design patterns (Alexander, 1977; Gamma et al., 1994) used in architecture and software engineering. For each, we define the strategy, explain the effects on objectives that motivate using the strategy, and how to implement the strategy with tactics. We then discuss secondary benefits, tradeoffs, examples from the literature, and connections to other strategies. While we note some differences between the crowdsourcing and chaining literatures throughout this section, we leave the bulk of such discussion to Section 7.
5.1. Diverse responses
To improve accuracy or creativity, source multiple responses to the same subtask with parallel, iterative, or debate based workflows (Figure 5).
Motivation. Individual workers can give low quality outputs, affecting the quality of the overall output. Designers may want to improve robustness to low quality outputs, outliers, and spammers to support the accuracy of outcomes (Liem et al., 2011; Du et al., 2023; Zaidan and Callison-Burch, 2011; Huang et al., 2018; Lasecki et al., 2013; Huffaker et al., 2020; Reppert et al., 2023; Ambati et al., 2012; Hahn et al., 2016). For creative tasks, they may want to encourage divergent and diverse thinking (Liang et al., 2023; Wu et al., 2023a; Luther et al., 2015; Lasecki et al., 2015; Huang et al., 2017; Nebeling et al., 2016; Teevan et al., 2016).
Implementation. Getting responses from multiple workers on a subtask can help support these goals. Crowdsourcing workflows and LLM chaining source diverse responses differently. Crowdsourcing does so by querying different people. Chaining leverages the stochastic nature of LLMs (e.g., with a higher temperature parameter setting) or or uses different models or prompt variants (including role-based prompts) to source a variety of responses (Park et al., 2022; Wu et al., 2023c; Arora et al., 2022). In some cases one can automatically calculate the optimal mix of LLMs for a task (Madaan et al., 2023a).
Redundant architectures are the most direct way of sourcing diverse responses, but communicative architectures, such as those that structure debates, can source diverse responses as well. Some studies in crowdsourcing systematically study the impact of different architectures (e.g., parallel vs iterative) and architecture parameters (e.g., the number of iterations) on the diversity and quality of the outcomes (Goto et al., 2016; Little et al., 2010a; Kittur et al., 2011; André et al., 2014). However, it is unknown if these findings transfer to similar tactic choices with LLMs.
Secondary benefits. There are some benefits of diverse responses that are solely talked about in the crowdsourcing literature, as they are dependent on having different people. Crowdsourcing brings up the benefit of having different viewpoints (Chen et al., 2019; Ambati et al., 2012; Lasecki et al., 2013), filling in each other’s deficits to create a more rounded final product (Kim and Monroy-Hernandez, 2016; Zhang et al., 2012), and sourcing perspectives from expert crowdworkers (Chilton et al., 2014) or from workers with various skill levels (Mahyar et al., 2018).
Consequences. Sourcing diverse responses is inherently more costly because multiple workers complete the same subtask. Merging multiple responses can induce outcome quality errors: coherency for parallel architectures (Wu et al., 2022a) and cascading errors for iterative architectures (Goto et al., 2016).
Examples. Examples of the diverse responses strategy include: Apparition (Lasecki et al., 2015), chains for multi-agent debate (Liang et al., 2023), and this exploration of ”human computation processes” (Little et al., 2010a).
Related strategies. Agreement among diverse responses can serve as validation. Setting quality thresholds can impact the number of diverse responses requested.
Title. Diverse responses, to improve accuracy or creativity, source multiple responses to the same subtask with parallel, iterative, or debate based workflows. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses, Coherency and cascading errors), Constraints (increase expense and latency), Accuracy (Robust to inaccurate outputs), Creativity (Improve with divergent thinking). Section 2: tactics used. Workers (LLMs, crowdworkers), Subtasks (all), Architectures (redundant, communicative).
5.2. Validation
To improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or worker-based checks (Figure 6).
Title. Validation, to improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or worker-based checks. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses), Constraints (increase expense and latency), Accuracy (check for correctness), Creativity (confirm elements of quality exist), Task specific (can validate specific qualities). Section 2: tactics used. Workers (LLMs, crowdworkers, user), Subtasks (evaluate, transform), Architectures (sequential, redundant, communicative).
Motivation. Individual workers can give low quality responses. Unaddressed, these can lead to cascading errors through the workflow (Wu et al., 2022a; Goto et al., 2016; Little et al., 2010b; Bernstein et al., 2010; Little et al., 2010a). Therefore, higher quality subtask outcomes throughout the workflow can have myriad benefits on the overall outcome quality for both accuracy and creativity focused tasks, increasing robustness to low-quality worker responses (Ambati et al., 2012; Bernstein et al., 2010).
Implementation. The validation process is often sequential, in which content is first generated before being validated (Gero et al., 2023). In parallel and iterative redundant workflows, errors from prior passes can be found and fixed or rejected (Zaidan and Callison-Burch, 2011; Hahn et al., 2016) and differing opinions can be merged (Kittur et al., 2011; Luther et al., 2015) through transform and evaluate subtasks (Wu et al., 2023a, c; Zharikova et al., 2023; Kulkarni et al., 2012; Ambati et al., 2012; Schick et al., 2022). Users can break ties (Pietrowicz et al., 2013) and intervene if there are cascading errors (Kulkarni et al., 2012). Non-worker methods, such as taking the majority vote of the outputs of redundant evaluate subtasks can also be effective. Automated methods of flagging malicious workers, gold-standard validation steps, or requiring a certain agreement threshold among workers within a redundant architecture can also provide validation (Lasecki et al., 2015; Parameswaran et al., 2023; Huffaker et al., 2020; Liem et al., 2011).
Debate-based communicative workflows can promote reflection on disagreements (Kobayashi et al., 2018; Chen et al., 2019; Du et al., 2023; Liang et al., 2023). Using different models and roles in debate shows promise (Du et al., 2023), as an LLM judge in a debate was found to more likely agree with its same underlying model (Liang et al., 2023). Other workflows use adaptable architectures to make all production always editable so workers can fix errors in a adaptable manner (Liu et al., 2018; LaToza et al., 2014).
Consequences. Costs increase when validation involves more calls to workers. The literature shows mixed results for an LLM’s ability to validate its own responses in evaluate subtasks. Some studies suggest that adding a self-validation step improves outcomes (Madaan et al., 2023b; Xie et al., 2023; Gero et al., 2023; Yao et al., 2023), while others suggest that LLMs are likely to agree with their previous outputs, even when incorrect, and propose debate-based communicative workflows as an alternative (Liang et al., 2023; Du et al., 2023).
Examples. Examples of the validation strategy include: the “Verify” step of Soylent (Bernstein et al., 2010), a Dual-Pathways structure for transcription (Liem et al., 2011), and self-evaluation guided reasoning (Xie et al., 2023).
Related strategies. Validation often involves considering the agreement among diverse responses.
5.3. User guidance
To help the user explore the design space and align creative outcomes to their vision, incorporate their input into the workflow at runtime (Figure 7).
Title. User Guidance. To help the user explore the design space and align creative outcomes to their vision, incorporate their input into the workflow at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Users learn through interaction), Constraints (Increased effort), Accuracy (Review uncertainties), Creativity (Guide the creative vision), Task general (Can predefine less before runtime). Section 2: tactics used. Workers (user), Subtasks (evaluate, focus), Architectures (redundant, dynamic, communicative).
Motivation. Users may want to use the workflow to reduce the effort of exploring of the design space, to align the workflow’s output to their creative vision, or some combination thereof (Nebeling et al., 2016; Willett et al., 2012; Kulkarni et al., 2014; Park et al., 2022). If a process includes data that is inherently uncertain it could benefit from more explicit guidance by the user (Chang et al., 2017; Willett et al., 2012; Cheng and Bernstein, 2015).
Implementation. This strategy incorporates the user into a critical role within the workflow. Users can guide outcomes by dynamically answering questions and providing feedback (Salehi et al., 2017) or monitor dynamic workflows to catch cascading errors (Kulkarni et al., 2012) with evaluate or transform subtasks. Users can be called upon when necessary with a “panic button” that other workers can invoke (Kulkarni et al., 2012), or automatically when ties occur (Pietrowicz et al., 2013). With a shared communicative interface between users and all workers, users can intervene when needed and make real-time suggestions for prototyping directions (Nebeling et al., 2016; Lasecki et al., 2015). Users may review at intermediate steps to edit or regenerate sections of the workflow in evaluate or focus subtasks (Park et al., 2022; Kim et al., 2023b; Zhang et al., 2023a; Kulkarni et al., 2014). A redundant iterative architecture may also include multiple iterations with user revisions between each one (Willett et al., 2012; Cheng and Bernstein, 2015; Kulkarni et al., 2014; Park et al., 2022; Kim et al., 2023b). Sometimes disagreements raised with redundant architectures may not need to be resolved within the workflow and can instead support an explorable interface for users to directly manipulate (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019). They can sort and filter the returned data (Willett et al., 2012; Cheng and Bernstein, 2015; Mohanty et al., 2019), use a slider to control output characteristics like text length (Bernstein et al., 2010), and consider category labels at multiple levels of precision (Chang et al., 2017).
Crowdsourcing workflows integrate the user in more varied ways than LLM chaining, including providing results without aggregation that the user can sort and filter (Lin et al., 2014a; Mohanty et al., 2019), user guidance for prototyping (Cheng and Bernstein, 2015; Kulkarni et al., 2014; Retelny et al., 2014), and worker-initiated user revision (Salehi et al., 2017).
Secondary benefits. Incorporating the user into the process at runtime can reduce the amount of pre-definition of workflow steps needed, making workflows more generally applicable (Nebeling et al., 2016; Kulkarni et al., 2014).
Consequences. These benefits come at the cost of more effort for the user than an entirely automatic workflow.
Examples. Examples of the user guidance strategy include: Second Opinion (Mohanty et al., 2019), Social Simulacra (Park et al., 2022), and WearWrite (Nebeling et al., 2016).
Related strategies. Incorporating user guidance can aid with validation.
5.4. Adaptable architectures
To build workflows that require less effort to design and that generalize to many inputs, use adaptive architectures that build themselves at runtime (Figure 8).
Title. Adaptable Architectures. To build workflows that require less effort to design and that generalize to many inputs, use adaptive architectures that build themselves at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Cascading errors and infinite loops, Less transparency and consistency), Constraints (Reduce expense and latency), Creativity (Adapt to edge cases), Task general (Adaptable for task generality, Request tasks with no pre-definition, Decomposable input required). Section 2: tactics used. Subtasks (focus, partition), Architectures (branching, dynamic, communicative).
Motivation. It is challenging and requires high effort to design a rigid and pre-defined workflow that can effectively generalize to many tasks or that can accommodate for many edge-cases, especially in creative tasks (Kim and Monroy-Hernandez, 2016; Valentine et al., 2017; Kittur et al., 2011; Kulkarni et al., 2012). Instead, designers may want workflows that are adaptable and responsive (Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2014; Chen et al., 2019; Wu et al., 2023a; Nair et al., 2023b; Creswell and Shanahan, 2022) and to support a variety of inputs without the effort of explicit pre-definition (Kittur et al., 2011; Kulkarni et al., 2012; Creswell et al., 2022; Creswell and Shanahan, 2022).
Implementation. Adaptable workflows are not rigid in their progression of subtasks but instead are flexible and generate parts of the workflow architecture at runtime (Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2014; Chen et al., 2019; Kittur et al., 2011; Huang et al., 2023c, b). Workflows can incorporate subtasks defined by the crowd with partition and focus subtasks in dynamic and communicative architectures (Ahmad et al., 2011; Zhang et al., 2012; Agapie et al., 2015; Retelny et al., 2014; Reppert et al., 2023). Changes in these adaptable workflows may be made by the worker (LaToza et al., 2014; Valentine et al., 2017; Retelny et al., 2014), user (Retelny et al., 2014; Nebeling et al., 2016; Valentine et al., 2017; Kulkarni et al., 2012), or an algorithm (Bragg et al., 2013; Chen et al., 2019; Dai et al., 2013; Goto et al., 2016). Although branching architectures are predefined, they also allow for adaptability in which different inputs take different paths.
Consequences. However, there are tradeoffs. Adaptable workflows can lead to poor quality worker-created subtasks and instructions, causing cascading errors and infinite loops (Kulkarni et al., 2012). They also are restricted to tasks with an easily decomposable input (Kittur et al., 2011).
Chaining papers highly emphasize the importance of transparency (Creswell et al., 2022; Creswell and Shanahan, 2022; Reppert et al., 2023; Wu et al., 2022a; Reppert et al., 2023; Gero et al., 2023; Wu et al., 2021, 2023a), increasing the risk of using adaptable workflows which may cause unforeseeable errors as each step is not predefined. The degree to which adaptable workflows degrade this transparency is yet to be explicitly explored.
Some effects of adaptable workflows on worker experiences mattered for crowdsourcing, but do not matter for chaining. Crowdsourcing work has allowed workers to chose their own tasks as an incentive to improve task performance (Chilton et al., 2019; Nebeling et al., 2016; Mahyar et al., 2018; Zhang et al., 2012; Kittur et al., 2011; LaToza et al., 2014), but the need to incentivize workers does not apply to chaining. One crowdsourcing study that found users were frustrated with too much adaptability, an emotion that does not apply to LLMs (Valentine et al., 2017).
Title. Worker recruitment, To improve outcome quality at each step, curate a high quality worker pool by filtering workers by their skills and improving worker capabilities over time. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increase cost and effort of filtering). Section 2: tactics used. Workers (LLMs, crowdworkers).
Examples. Examples of adaptable architectures include: map-reduce (Kittur et al., 2011), Flash Teams (Retelny et al., 2014), and microtask programming (LaToza et al., 2014).
Related strategies. Adaptable architectures must support the dynamic creation of aligned subtasks that have well defined context and worker instructions. Adaptable architectures can also dynamically change course once quality thresholds are met.
5.5. Worker recruitment
To improve outcome quality at each step, curate a high quality worker pool by filtering workers by their skills and improving worker capabilities over time (Figure 9).
Motivation. As workers power the workflow, workers that can provide higher-quality responses can help support outcome quality (Bragg et al., 2018; Liu et al., 2016).
Implementation. In crowdsourcing, recruiting high-quality workers can mean filtering crowdworkers by pre-specified qualifications or using platforms with more specialized workers such as Upwork (Valentine et al., 2017). For LLM chains, model choice and finetuning impact worker abilities (Cobbe et al., 2021; Bursztyn et al., 2022; Creswell et al., 2022; Creswell and Shanahan, 2022).
Consequences. Higher quality workers may be more expensive (Christoforaki and Ipeirotis, 2014; Huang et al., 2015; Bursztyn et al., 2022). There is also effort and expense involved in filtering crowdworkers for skill (Kulkarni et al., 2014; Zaidan and Callison-Burch, 2011; Valentine et al., 2017) or finetuning models to improve specific skills (Bursztyn et al., 2022).
Examples. Examples that pay particular attention to worker recruitment include Flash Organizations (Valentine et al., 2017), and chains that finetune models (Bursztyn et al., 2022) or use multiple models (Liang et al., 2023).
Related strategies. Worker recruitment works hand in hand with aligning subtasks to worker capabilities. Worker recruitment must also be considered when using adaptive architectures to define what kind of workers to hire during runtime.
5.6. Subtask alignment
To improve outcome quality at each step, align subtasks to worker capabilities (Figure 10).
Title. Subtask alignment, To improve outcome quality at each step, align subtasks to worker capabilities. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increased effort to design). Section 2: tactics used. Subtasks (partition, focus), Architectures (dynamic, communicative).
Motivation. One critical strategy is ensuring that subtasks match the workers’ capabilities appropriately. Simplifying tasks can increase outcome quality (Chilton et al., 2013; Wu et al., 2022b; Lasecki et al., 2013; Pietrowicz et al., 2013).
Implementation. Compared to entire tasks, smaller subtasks are often easier to complete, validate, and correct (Ambati et al., 2012; Schick et al., 2022; Wu et al., 2022b; Yang et al., 2022; Kim et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022). Designers can also adapt the subtask content to fit the worker’s capabilities (Chang et al., 2017; Mohanty et al., 2019; Pietrowicz et al., 2013; Arora et al., 2022). For example, researchers found that crowdworkers were better at generating predictive features than at estimating if a feature is predictive; so they adapted the workflow accordingly (Cheng and Bernstein, 2015).
When the appropriate subtasks are unknown, dynamic architectures can use partition or focus subtasks to use workers’ input to break down tasks into an appropriate size at runtime (Kittur et al., 2011; LaToza et al., 2014; Kim et al., 2017; Huang et al., 2023c). Communicative architectures can create appropriate subtasks for workers when they have a running “todo” list from which workers can choose their preferred task of many (Zhang et al., 2012; Lasecki et al., 2015; Mahyar et al., 2018).
Consequences. Much remains unknown about creating appropriate subtasks for LLMs. The capabilities of LLMs and appropriate subtask content are far from fully understood. Furthermore, it is unknown the degree to which LLMs can judge the appropriateness of subtasks for other LLMs to complete, bringing into question the effectiveness of some partition subtask and communicative architecture approaches taken when the workers are crowdworkers. Therefore, testing out these different options to find appropriate subtasks is resource intensive in the design process.
Examples. Examples that pay particular attention to subtask alignment include: Revolt (Chang et al., 2017), CrowdBand (Pietrowicz et al., 2013), and Ask Me Anything (Arora et al., 2022).
Related strategies. Subtasks must align to the capabilities of recruited workers and requirements must be communicated with sufficient context. Subtask alignment is especially challenging when subtasks are created during runtime eith adaptable architectures.
5.7. Context
To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of worker instructions (Figure 11).
Title. Context, To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of worker instructions. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Critical for quality outputs, Excessive context can confuse), Constraints (Increase effort to design), Accuracy (Affected by small changes). Section 2: tactics used. Subtasks (all), Architectures (communicative).
Motivation. Workflows need to have high quality subtask outputs that promote cohesion among disparate parts of the workflow.
Implementation. Communicating context to workers can help with improving the quality of the output and creating more coherence among disparate subtasks (Luther et al., 2015; Wu et al., 2022a; Teevan et al., 2016; Chilton et al., 2019; Hahn et al., 2016; Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010). There are many forms of contextual communication, from input text (Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010), to workflow goal tracking (Mirowski et al., 2023; Yang et al., 2022; Kim et al., 2017), to explicit communication between users and workers (Salehi et al., 2017).
Instruction wording is a critical part of all subtask designs to support high outcome quality (Parameswaran et al., 2023). A key difficulty is communicating the user’s intent through language. One study reports that crowdworkers tend to need less explicitly defined instruction than LLMs, but can still struggle with vague or ambiguous instructions (Wu et al., 2023c). Both crowdworkers and LLMs benefit from examples, namely screening and tutorial tasks for crowdworkers (Teevan et al., 2016; Lasecki et al., 2015; Chen et al., 2019; Lasecki et al., 2013; Pietrowicz et al., 2013; Drapeau et al., 2016; Bernstein et al., 2010) and few-shot examples for LLMs (Nair et al., 2023a; Gero et al., 2023). Interfaces for crowdworkers can structure their output, while this structure needs to be communicated to LLMs in natural language (Wu et al., 2023c; Creswell and Shanahan, 2022; Kim et al., 2014).
Consequences. Although context is generally important, overcomplicating instructions can lead to confusion (Chen et al., 2019; Lasecki et al., 2015; Wu et al., 2023c; Madaan et al., 2023b), so understanding the right amount of context to give is important (Kittur et al., 2011). One in-depth crowdsourcing study on providing context through interactions with the user found that providing context was helpful early on in the workflow if content is poor, but became detrimental if communicated later or with high-quality content (Salehi et al., 2017). Minimizing context seen can also help with preserving privacy (Bragg et al., 2021; Kaur et al., 2017). For LLMs, context may be included as part of the prompt, although LLMs have the additional barrier of an input token limit (Park et al., 2022; Mirowski et al., 2023). Therefore, workflow designs should aim for a minimum amount of sufficient context, to provide coherence while avoiding confusion.
[…]